OpenStudy (anonymous):

A ball is thrown from an initial height of 2 feet with an initial upward velocity of 21f/s. The ball's height (in feet) after seconds is given by the following. h= 2 + 21t - 16t^2 Find all values of T for which the ball's height is 8 feet. Round your answer(s) to the nearest hundredth.

OpenStudy (irishboy123):

h(T) = 2 + 21T - 16T^2 = 8 solve 16T^2 - 21T + 6 = 0 for T